is 8 * c * 3
8c * 3
or
40c?
why?

Answer by jim_thompson5910(35256)

8*c*3 = 8*3*c = 24c

So 8*c*3 = 24c

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------